OpenStudy (anonymous):

May Somebody please Help Me ... The numbers of seats in the first 12 rows of a high-school auditorium form an arithmetic sequence. The first row has 9 seats. The second row has 11 seats. A. Write a recursive formula to represent the sequence. B. Write an explicit formula to represent the sequence. C. How many seats are in the 12th row?

OpenStudy (smileyxl3):

a) R₁ = 9 R(n) = R(n−1) + 2 b) R(n) = 9 + 2 * (n−1) R(n) = 7 + 2n c) R₁₂ = 7 + 24 = 31

OpenStudy (anonymous):

@smileyxl3 thank you SOOO much you are a life savior

OpenStudy (smileyxl3):

@xoxobee You're Welcome